OpenStudy (anonymous):

magine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

First I would tell the student to use the distributive property on everything inside the parenthesis 4-x-2=6-5x-20

OpenStudy (anonymous):

Okay and after what would you do ... ?

OpenStudy (radar):

1. Remove the parenthesis by distributing the -1 (on the left) and the -5 (on the right) this will give you: 4-x-2=6-5x-20 2. Add 5x to both sides of the equal sign in order to get the x terms on the left side. this will give you: 4-4x-2=6-20 3. combine constants giving you: -4x+2=-14 4. Subtract 2 from both sides to get constant on the right. Giving you: -4x=-16 5. Divide both sides by -4 giving you: x=4

OpenStudy (anonymous):

Thanks alot Radar.

OpenStudy (radar):

I hoped it helped, There are various ways this could of been done. I showed you the way I would of done it. You are welcome.